Gucchi:

Which of the following equations represents the greatest value of x? Equation 1 Equation 2 Equation 3 Equation 4 x2 = 24 x3 = 64 x2 = 28 x3 = 125 Equation 1 Equation 2 Equation 3 Equation 4

Gucchi:

Idk if its c or d

Ashoka:

What I would do here is to solve for x, for example x^3=125. then x = 5

Ashoka:

Because 5 times 5 times 5 = 125

darkknight:

No direct answers, Adding onto Ashoka, You want to solve for x, so for x, x^3=64, then x = 4 because 4 times 4 times 4 = 64

Gucchi:

Ashoka thats true

Gucchi:

What do i do after i figure out x?

Gucchi:

for each one

darkknight:

Compare them in each term, so once you find x for each one see which one is the greatest

Gucchi:

see which x is the greatest?

darkknight:

Yes

Gucchi:

Oh ok

Gucchi:

i can't figure out what goes into 24 and 28 two times

darkknight:

For that square root both sides, so that x^2=38 \[x = \sqrt(28)\]

darkknight:

sorry x^2=28 not 38, my b

Gucchi:

Uhm

Gucchi:

I'm so confused

darkknight:

Take a calculator, what is the square root of 28???

Gucchi:

5.29150262213

Gucchi:

If i go on brainly, why are people dividing numbers?

darkknight:

Your job is to compare x values, and don't ask about brainly, idk and idc. You have a squared term so you have to square root it, so which equation did you find x to be the greatest in?

Gucchi:

equation three

Ashoka:

That is correct,

darkknight:

Yep correct, have a good day : )

Gucchi:

it has the greatest

Gucchi:

Thanks

darkknight:

np